OpenStudy (anonymous):

Use long division to determine the decimal equivalent of fraction with numerator 5 and denominator 9. 0 point 3 bar 0 point 4 bar 0 point 5 bar 0 point 6 bar

OpenStudy (anonymous):

0.5 Bar

OpenStudy (anonymous):

I can has medal plox?

OpenStudy (anonymous):

thx :D

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

Btw

OpenStudy (anonymous):

ur welcome :D

OpenStudy (anonymous):

WHenever a number is over 9, or you can simplify or complicate it to 9, Say 5/9 itll always be 0.(number) in this case 5 bar

OpenStudy (anonymous):

ok, can u help me some more???

OpenStudy (anonymous):

??? @ethanxiaorox

OpenStudy (anonymous):

Betty had $70. She bought a necklace for $9 and a ring for $7. Which expression correctly shows the total money Betty has left? A $70 − (−$9) + (−$7) B $70 + (−$9) + (−$7) C $70 − (−$9) − (−$7) D $70 + (−$9) − (−$7) A B C D

OpenStudy (anonymous):

B

OpenStudy (anonymous):

But put it this way.

OpenStudy (anonymous):

1/9=1 bar, 2/9 = 2 bar, Etc

OpenStudy (anonymous):

Note that only works with 1 digit numbers, for 2 digit, you do 10/99 = 10 bar, 11/ 99 =11 bar, 12/99 = 12 bar Etc

OpenStudy (anonymous):

thx and...

OpenStudy (anonymous):

In a music competition, a participant has to score a total of at least 40 points in the first four rounds combined to move on to the fifth and final round. Glenn scored 4 points in the first round. He then went on to score additional points in the second, third, and fourth rounds. In each of those rounds, his score was identical. Which inequality best shows the number of points, p, that Glenn scored in each of the second, third, and fourth rounds if he earned a place in the finals? 4 + 3p ≤ 40 4 + 3p ≥ 40 4p + 3 ≥ 40 4p + 3 ≤ 40

OpenStudy (anonymous):

\[the first one is \le second \ge third \ge fourth\]

OpenStudy (anonymous):

\[fourth is \]

OpenStudy (anonymous):

B

OpenStudy (anonymous):

the fourth one is like the first one thx :D